OpenStudy (anonymous):

Use the law of Cosines to solve the triangle. B=12º, a=32, c=36

OpenStudy (psymon):

So we just need to plug some numbers into the equation for the law of cosines and go from there. Now I can choose to rearrange the formula in different ways as long as you do it properly. So I'll list the formula like this and see if you can plug the appropriate numbers in and solve for the angle

OpenStudy (anonymous):

ok

OpenStudy (psymon):

\[a ^{2} = b ^{2} + c ^{2} - 2ac(\cos A)\]

OpenStudy (anonymous):

32^2= b^2=36^2

OpenStudy (anonymous):

i mean + after b^2

OpenStudy (anonymous):

but i dont have cos A

OpenStudy (psymon):

You need to solve for cos A. You can fill in all the information and eventually be left with something = cosA

OpenStudy (psymon):

\[32^{2} = 12^{2} + 36^{2} -2(12)(36)(cosA)\] First solve for cosA = something and we'll go from there :3

OpenStudy (anonymous):

is it 448?

OpenStudy (psymon):

If done corrwctly, you should get \[\frac{ 13 }{ 27 } \]

OpenStudy (anonymous):

as the final answer?

OpenStudy (psymon):

Im sorry, it got cut off, it shouldve said 13/27 = cosA

OpenStudy (psymon):

Do you know how to solve for cosA from here?

OpenStudy (anonymous):

what are they exactly asking to do ?

OpenStudy (psymon):

They want you to find angles A, B, and C. When they give you lower case letters, these refer to the sides of a triangle. Upper case letters refer to angles. So solving a triangle is having all 3 sides and all 3 angles. In this case you had all 3 sides but no angles. so we're trying to find the 3 angles. So right now, we need A. We've made it up to cosA = 13/27, now we just need to get A. Any idea how to do that?

OpenStudy (anonymous):

do we use the inverse of cosine?

OpenStudy (psymon):

Exactly.

OpenStudy (anonymous):

is it tan^-1 pop/adj?

OpenStudy (anonymous):

i mean opp srry

OpenStudy (psymon):

Well, this is a calculator question really. You would need to do: \[\cos^{-1} (\frac{ 13 }{ 27 }) = \theta \]

OpenStudy (anonymous):

oh ok got it

OpenStudy (psymon):

Yep. So what'd we getfor theta?

OpenStudy (anonymous):

whats the formula for the law of sines? 61.21

OpenStudy (psymon):

Sorry about that. And yes, 61.21. From there, the formula is: \[\frac{ a }{ sinA } = \frac{ b }{ sinB } = \frac{ c }{ sinC }\]

OpenStudy (anonymous):

how do we proceed after this?

OpenStudy (psymon):

Well, we have an angle A and a side a, so this allows us to use this information to solve for either of the other two angles. If I wanted to solve for B first, I would do: \[\frac{ 32 }{ \sin(61.21) } = \frac{ 12 }{ sinB }\] from here we can solve for sinB and, once again, take the inverse once we finish.

OpenStudy (anonymous):

how doe we solve for sin B?

OpenStudy (psymon):

If we have equivalent fractions then this means we can cross multiply and eventually get sinB by itself.

OpenStudy (anonymous):

got it